OpenStudy (anonymous):

Input: 0 1 2 3 Output: 5.5 4.2 3.2 2.2 What's the relationship ?

OpenStudy (anonymous):

@AccessDenied can u help me please ?

OpenStudy (anonymous):

when Input increase, output will decrease.

OpenStudy (nubeer):

ya seems like linearly decreasing.

OpenStudy (anonymous):

it is not linear..

OpenStudy (anonymous):

output=-1.3(input)+5.5

OpenStudy (anonymous):

linear decreasing

OpenStudy (anonymous):

yes

OpenStudy (nubeer):

@cinar it looks linear decreasing , if one value is a lil diff doesnt mean its not linear.

OpenStudy (anonymous):

no for first two values

OpenStudy (anonymous):

There is no formula available. It is not an arithmetric progression because the distance between values is not constant It is not a geometric progression between the ratio of consecutive values is not constant

OpenStudy (anonymous):

y=-0.05x^3+0.3x^2-1.55x+5.5

OpenStudy (nubeer):

it's a smiple question .. why making so much complicated.

OpenStudy (nubeer):

why .. doesn't it look linear?

OpenStudy (anonymous):

assume that f(x)=ax^3+bx^2+cx+d f(0)=5.5=d f(1)=a+b+c+5.5=4.2 f(2)=8a+4b+2c+5.5=3.2 f(3)=27a+9b+3c+5.5=2.2

OpenStudy (anonymous):

a+b+c=-1.3 8a+4b+2c=-2.3 27a+9b+3c=-3.3 after solve this eq. system you'll get a,b,c coeff.

OpenStudy (anonymous):